OpenStudy (anonymous):

\(\Large \int\limits \frac{ (1+e^x)^2 }{ e^x } dx\) Prove by induction that for \( x \ge 0 \) and any natural number \(n\), \(\large e^x \ge 1+x+\frac{ x^2 }{ 2! } + ... + \frac{ x^n }{ n! }\)

OpenStudy (anonymous):

two problems.

OpenStudy (loser66):

for the first one, to me, I break it down to (1+2e^x +e^x^2)/e^x= e^(-x) +2 +e^x, then take integral term by term.

OpenStudy (anonymous):

yes,that idea does word :/ ... and for second one?

OpenStudy (zarkon):

for the 2nd one 1) do the basis step 2) assume \(\large e^x \ge 1+x+\frac{ x^2 }{ 2! } + ... + \frac{ x^n }{ n! }\) is true for some \(n\) let \(\displaystyle f(x)= e^x -\left(1+x+\frac{ x^2 }{ 2! } + ... + \frac{ x^n }{ n! }+\frac{ x^{n+1} }{ (n+1)! }\right)\) show that \(f(x)\) is an increasing function (by taking the derivative) that is only zero when \(x=0\)